### v2

# High-Fidelity Simultaneous Speech-To-Speech Translation

Abstract
We introduce Hibiki, a decoder-only model for simultaneous speech translation. Hibiki leverages a multistream language model to synchronously process source and target speech, and jointly produces text and audio tokens to perform speech-to-text and speech-to-speech translation. We furthermore address the fundamental challenge of simultaneous interpretation, which unlike its consecutive counterpart—where one waits for the end of the source utterance to start translating— adapts its flow to accumulate just enough context to produce a correct translation in real-time, chunk by chunk. To do so, we introduce a weakly-supervised method that leverages the perplexity of an off-the-shelf text translation system to identify optimal delays on a per-word basis and create aligned synthetic data. After supervised training, Hibiki performs adaptive, simultaneous speech translation with vanilla temperature sampling. On a French-English simultaneous speech translation task, Hibiki demonstrates state-of-the-art performance in translation quality, speaker fidelity and naturalness. Moreover, the simplicity of its inference process makes it compatible with batched translation and even real-time on-device deployment. We provide examples [^1] https://hf.co/spaces/kyutai/hibiki-samples as well as models and inference code. [^2] https://github.com/kyutai-labs/hibiki
Machine Learning, ICML
1 Introduction
We introduce Hibiki (“echo” in Japanese), a system for streaming and expressive speech-to-speech (S2ST) and speech-to-text (S2TT) translation. Most work in speech translation has focused on the offline (or consecutive) setting where the model has access to the full source utterance before it translates, as it provides useful context while fitting many use cases such as offline video dubbing. A more challenging setting is that of simultaneous translation, where translated speech is produced on-the-fly. This task, typically performed by human interpreters, requires a real-time decision-making process to evaluate whether enough context has been accumulated to reliably translate another chunk of speech. When cast as a machine learning problem, this endeavor presents additional challenges such as the lack of speech data aligned at a chunk-level.
Hibiki is a decoder-only model which synchronously receives source speech and generates translated speech by leveraging a multistream architecture. In this context, nested global and local Transformers (Vaswani et al., 2017) jointly model two audio streams by predicting a hierarchy of text and audio tokens for each of them. At inference time, temperature sampling combined with a causal audio codec allows for streaming inputs and outputs. While this architecture was originally introduced by Défossez et al. (2024) for full-duplex spoken dialogue, we show how it provides a simple and convenient architecture for simultaneous speech translation. To train Hibiki, we generate synthetic parallel data by translating and resynthesizing the transcript of single-language audio. While this provides pairs of inputs and outputs aligned at the sequence level, this does not allow learning fine-grained alignments. We thus introduce “contextual alignment”, a simple method based on the perplexity of an off-the-shelf machine translation system (Kudugunta et al., 2023) to derive word-level alignments. By then introducing proper silences into target speech, we can train Hibiki to adapt its flow in real-time, without the need for complex inference policies. Moreover, observing that training data varies widely in speaker similarity, we propose to label training examples with categories of speaker similarity, which avoids filtering the training data while allowing to favor high speaker similarity at inference with classifier-free guidance.
In a French-English translation task, Hibiki outperforms previous work in translation quality, speaker similarity and naturalness. We also show how the simplicity of our inference process allows for translating hundreds of sequences in real-time on GPU, and how a distilled model can run in real-time on a smartphone. Human evaluations demonstrate that Hibiki is the first model to provide an experience of interpretation close to human professionals. We will release our code and models, and a high quality 900 hours synthetic dataset.
2 Related Work
Figure 1: Generating aligned interpretations. We extract unsupervised word level contextual alignment, which we lift to audio by either inserting silences, or re-synthesising with an alignment aware TTS. See Section 3.2 for details.
2.1 End-to-end speech translation
Speech translation can be traced back to the early 1990s (Jain et al., 1991) with a first generation of systems that combined automatic speech recognition (ASR), machine translation (MT) and text-to-speech synthesis (TTS). While such cascaded approaches allowed for the growth of speech translation (Wahlster, 2000; Nakamura et al., 2006), they suffer from two main limitations. First, they are subject to compounding errors due to combining separately trained models. This motivated the merging of ASR and MT into a single speech-to-text translation (S2TT) model (Berard et al., 2016; Weiss et al., 2017; Wang et al., 2020, 2021b) that can provide inputs to a TTS model. However, a second limitation of cascaded systems remains: as the input speech goes through a text bottleneck, the non-linguistic information it carries—such as speaker identity or prosody—is lost and cannot be transferred to the output speech. End-to-end speech-to-speech translation (S2ST) (Jia et al., 2019; Lee et al., 2022a; Jia et al., 2022a; Rubenstein et al., 2023) addresses this issue by directly predicting target speech from source speech, allowing for retaining paralinguistic information, including voice identity. A notable aspect of most end-to-end S2ST models is that they leverage auxiliary text or phoneme translation tasks in training, that are then discarded (Jia et al., 2022a) or run in parallel (Zhang et al., 2024a) to the main speech translation task at inference. Hibiki performs end-to-end S2ST with voice transfer along with S2TT, but instead of running these tasks in parallel, Hibiki uses the predicted text as a scaffolding for speech generation at inference time. Moreover, since Hibiki predicts aligned speech and text tokens, it provides word-level timestamps in the target language.
2.2 Simultaneous speech translation
While the first attempts at simultaneous speech translation focused on speech-to-text (Ren et al., 2020; Ma et al., 2021; Zeng et al., 2021), Seamless (Barrault et al., 2023) and StreamSpeech (Zhang et al., 2024a) have introduced end-to-end simultaneous S2ST (Zeng et al., 2021; Barrault et al., 2023; Zhang et al., 2024a). Both systems predict discrete speech units with autoregressive models before decoding them to audio using a neural vocoder, and rely on a specific policy for inference. While StreamSpeech translates into a canonical voice, Seamless performs voice transfer from source to target. Hibiki also performs simultaneous S2ST and S2TT, while transferring voice characteristics. However, Hibiki relies on a decoder-only model which operates at a constant frame rate and performs inference with simple temperature sampling or greedy decoding. In particular, this allows for batching, unlike StreamSpeech’s and Seamless’s policies that involve a complex control flow that cannot be batched. This makes Hibiki able to translate hundreds of sequences on a single GPU, provides convenient support for classifier-free guidance, see e.g. Section 3.3, and allows to run in real time on device, as shown in Section 4.6.1. Human evaluations in Section 4.6 show that Hibiki significantly outperforms Seamless in terms of naturalness and audio quality, getting close to human interpretation.
3 Method
Figure 2: Joint sequence modeling with contextual alignment. From the source stream, Hibiki predicts its Inner Monologue text stream, and audio tokens. Its output is aligned for causality, as depicted in Figure 1. Figure adapted from  Défossez et al. (2024).
We consider an utterance in a source language represented as a monophonic waveform $X\in\mathbb{R}^{f_{s}\cdot d}$, sampled at a frame rate $f_{s}=24\,\mathrm{kHz}$, of duration $d$. Similarly, its translation is given in a target language, denoted $Y\in\mathbb{R}^{f_{s}\cdot d}$. We assume $X$ is padded to ensure both have the same duration. Our objective is to model $\mathbb{P}\left[Y|X\right]$. We further add the constraint that the modeling of $Y$ knowing $X$ should be causal and of minimal delay with respect to the source utterance, e.g. the same constraints that are imposed on a human interpreter in the context of live translation. To learn this constraint via supervised learning, $Y$ must itself be built to respect this causality constraint. We first assume that $Y$ respects this constraint, and we present how to model its distribution. Then, we introduce an information theory criterion to verify whether $Y$ is causal with respect to $X$, and to adapt a non-causal interpretation into a causal one.
3.1 Modeling
We build on the framework introduced by Défossez et al. (2024) for the joint modeling of multiple sequences of discrete tokens, obtained from a neural audio codec.
3.1.1 Neural audio codec
We use the pre-trained causal and streaming Mimi codec (Défossez et al., 2024) to encode $X$ and $Y$ into low framerate sequences of discrete tokens. Mimi consists of an encoder and decoder from and to the waveform domain, and of an information bottleneck using Residual Vector Quantization (RVQ) (Zeghidour et al., 2022). The encoder transforms an input waveform of duration $d$ into a latent vector $U\in\mathbb{R}^{C\times f_{r}\cdot d}$, with $C$ the dimension of the latent space, and $f_{r}=12.5\,\mathrm{Hz}$ the frame rate. $U$ is then projected to its nearest neighbor in a codebook table with $N_{A}$ entries. The residual of the projection is further projected into a second table with the same cardinality, and so forth until $Q$ projections have been performed. The last residual is discarded, and the decoder is trained to reconstruct the input waveform from the sum of the projected tensors. The codebooks are trained through exponential moving average, along with a commitment loss (Razavi et al., 2019). The rest of the model is trained only through an adversarial loss with feature matching (Défossez et al., 2024).
For language modeling, we are not interested in the quantized latent vector and its residuals, but in the discrete indices of the entry in the codebooks it is projected to. We denote those $(A_{t,q})\in\{1,\ldots,N_{A}\}^{f_{r}\cdot d\times Q}$. For Mimi we have $f_{r}=12.5\,\text{Hz}$ and $Q$ varies up to 32, but we use at most 16. Following Zhang et al. (2024b); Défossez et al. (2024), the output of the first quantization level is trained to replicate semantic information obtained from a WavLM self-supervised audio model (Chen et al., 2022). Following conventions of Borsos et al. (2022), we refer to $A_{t,1}$ as the semantic tokens, and $A_{t,q\geq 2}$ as the acoustic tokens. The acoustic tokens are arranged in a coarse to fine manner, the first ones have the most importance, and the latest model fine details of the audio and ensuring a smooth perception.
3.1.2 Joint modeling of discrete audio tokens
The discrete tokens for audio streams cannot easily be summarized into a single discrete sequence with reasonable cardinality and framerate (Copet et al., 2023). Following Yang et al. (2023); Défossez et al. (2024), we leverage a RQ-Transformer (Lee et al., 2022b) to model $(A_{t,q})$ both over the time $t$ and quantizer $q$ axes. It consists in a large Temporal Transformer (Vaswani et al., 2017), operating at the same framerate $f_{r}$ as the codec, and being fed all the tokens generated so far, e.g. for all $t\leq f_{r}\cdot d$,
$Z_{t}=\mathrm{Temp}(A_{0},\ldots,A_{t-1})\in\mathbb{R}^{D}.$ (1)
$A_{0}$ is defined as a deterministic token indicating the start of the generation. Then, a smaller scale Depth Transformer models auto-regressively the tokens $A_{t,1},\ldots,A_{t,Q}$ over the quantizer axis, e.g. for all $t\leq f_{r}\cdot d$ and $q\leq Q$,
$l_{t,q}=\mathrm{Dep}(Z_{t},A_{t,0},\ldots,A_{t,q-1})\in\mathbb{R}^{N_{a}},$ (2)
with $A_{t,0}$ also a special token, and with the goal of having,
$\mathrm{softmax}(l_{t,q})\approx\mathbb{P}\left[A_{t,q}|A_{0},\ldots,A_{t-1},A_{t,0},\ldots A_{t,q-1}\right]$
Following (Copet et al., 2023; Défossez et al., 2024), we further introduce an acoustic delay of 2 time steps, meaning that we model $(\tau(A)_{t,q})$ instead of $A$,
$\begin{cases}\begin{array}[]{lll}\tau(A)_{t,1}&=A_{t,1}&\quad\forall\,t\\ \tau(A)_{t,q}&=A_{t-2,q}&\quad\forall\,t\geq 3,\forall\,q\geq 2\\ \tau(A)_{t,q}&=0&\quad\forall\,t<3,\forall\,q\geq 2,\\ \end{array}\end{cases}$ (3)
with 0 being a special token. The delay is removed before decoding audio with the codec.
3.1.3 Translation as multistream modeling
We have presented how the RQ-Transformer given by eq. (1) and (2) allows for jointly modeling multiple discrete streams of tokens. We adapt this framework for the task of joint speech-to-speech and speech-to-text simultaneous translation. We concatenate the audio tokens $A^{Y}$ obtained from the target interpretation $Y$, with the tokens $A^{X}$ from the source utterance $X$ along the $q$-axis, e.g.
$\bar{A}=\mathrm{concat_{q}}\left[\tau(A^{Y}),\tau(A^{X})\right].$ (4)
We observe a benefit from modeling the tokens $A^{X}$ at train time, although at inference time, predictions for those tokens are skipped and actual tokens for the input are used instead.
Défossez et al. (2024) showed generating an Inner Monologue, i.e. padded text tokens aligned with the content of the generated audio, is beneficial to the quality and stability of the generated audio. This is similar to multi-task learning where the translation is predicted both in the audio and text domain. Hibiki thus also predicts a text stream corresponding to the transcription of the output $Y$, with sufficient padding between words to keep them aligned with the audio. Note that unlike previous multi-task translation work, Hibiki makes active use of this capability at inference time. We denote $W_{t}$ the text stream, with cardinality $N_{W}$ and the same frame rate $f_{r}$ as the audio streams.
3.1.4 Architectural details
We provide architectural hyper-parameters in Section 4.1. At time-step $t$, the tokens from the step $t-1$, e.g. $\tau(A^{X})_{t-1}$, $\tau(A^{Y})_{t-1}$, and $W_{t-1}$, are fed into dedicated embedding tables, whose contributions are summed. For the first time step $t=1$, a BOS token is used instead. We then use standard Transformer layers (Vaswani et al., 2017), with gated SiLU activation (Shazeer, 2020; Hendrycks & Gimpel, 2016). A linear layer maps its output $Z_{t}$ to logits for the text token $W_{t}$. The Depth Transformer then operates for $2\cdot Q$ steps: the first half to estimate the logits for the output stream, and the next half for the input stream. Each depth step $q$ takes as input $Z_{t}$ summed with a learnt embedding of the previous audio token $\bar{A}_{t,q-1}$, or $W_{t}$ for $q=1$.
3.2 Alignment and synthetic interpretation data
Figure 3: Contextual alignment. We compute the log-likelihood of the word “into” with a pre-trained text translation model, for various input truncations. Once the matching source word ”en” appears, we observe a large increase in log-likelihood, see eq. (6).
We have assumed pairs $(X,Y)$ that respect the constraint of simultaneous interpretation. We now introduce an unsupervised criterion to estimate and enforce causality dependencies between the source and target utterances.
3.2.1 Text domain alignments
Let us first express formally those constrained in the text domain. Let us take $S=(S_{1},\ldots,S_{n})$ the sequence of words in the utterance $X$, and $T=(T_{1},\ldots,T_{m})$ that in $Y$.
Ideal alignment.
We seek to define an ideal alignment $(a^{\text{ideal}}_{j})\in\{1,\ldots,n\}^{m}$, where $a^{\text{ideal}}_{j}$ indicates the index of the word in $S$ that the $j$-th word in $T$ should wait for to minimize the uncertainty on $T_{j}$. Any alignment strictly less conservative than $a^{\text{ideal}}$ would risk the model hallucinating at inference if trained on. Any alignment strictly more conservative would still be causal, but introduce more latency.
Contextual alignment.
We introduce a criterion to estimate $a^{\text{ideal}}$. Let’s denote the conditional log-likelihood
$\log(p_{j,i})=\log\left(\mathbb{P}\left[T_{j}|S_{1},\ldots S_{i},T_{1},\ldots,T_{j-1}\right]\right),$ (5)
we expect $\log p_{j,i}$ to increase with $i$, as more context is beneficial. We conjecture that $\delta_{j,i}=\log(p_{j,i})-\log(p_{j,i-1})$ is maximal for $i=a_{j}$. We compute an estimate $\log(\hat{p}_{j,i})$ of $\log(p_{j,i})$ with an off-the-shelf text translation language model MADLAD-3B (Kudugunta et al., 2023), by feeding it truncated input up to the $i$-th word, which we use to define a contextual alignment, illustrated in Figure 3,
$a^{\text{ctx}}_{j}=\operatorname*{arg\,max}_{i\leq n}\left[\log(\hat{p}_{j,i})-\log(\hat{p}_{j,i-1})\right].$ (6)
Examples of alignments are given in the Appendix, Figure 7.
3.2.2 Audio domain alignments
Given $(X,Y)$, we transcribe both with timestamps with a Whisper model (Radford et al., 2023; Louradour, 2023) and apply eq. (6). The pair $(X,Y)$ respects the alignment $(a^{\textrm{ctx}}_{j})$ if the timestamp of the $j$-th word in $Y$ comes after the timestamp of the $a_{j}$-th word in $X$. To reduce the impact of errors, we require $Y$ to lag by at least 2 seconds compared to the contextual alignment, and eliminate spikes higher than 25% of the average delay over a window of 5 words.
Silence insertion.
If $Y$ doesn’t respect the alignment, one can simply transform it by inserting sufficient silences before a word, as illustrated in Figure 1. This comes, however, with two limitations:
(i) silence insertion can lead to hard cuts when the timestamps are inaccurate or no pause exists between words;
(ii) the corrected $Y$ might be arbitrarily late on the ideal alignment, e.g. if the speech rate is slower in $Y$ than in $X$.
We apply this method during the speech translation training.
Alignment-aware TTS.
We obtain more natural aligned data by (re)-synthesising $Y$ with a TTS model able to follow hard and soft constraints on word locations, along with accurate speaker conditioning. For existing datasets, this can have the added benefit of improving the word error rate and the speaker similarity, as illustrated in Section 3.3. Following (Défossez et al., 2024), Appendix C, we train a TTS with both audio and a synced text stream as output, along with voice conditioning. The text stream is constrained to match exactly text to generate, with the model having only the freedom to insert padding tokens. The audio output is late on the text, so that its content is conditioned by it, both for content and timestamps. If the TTS is early on the alignment $a^{\textrm{ctx}}$, padding tokens are forced to delay the next word. When the TTS is lagging on its target, a penalty is added on the logits of the padding token. The penalty scales from 0 to -2 as the lag increases from 1 to 2 seconds. This increases smoothly the rate of speech to catch up with the source audio. We perform 6 to 8 generations per input, and select the best one based on word error rate first, and speaker similarity second. We apply this only for the fine-tuning speech translation dataset.
Figure 4: Speaker similarity between source and target speech in CVSS-T training data, before and after resynthesis.
3.3 Voice Transfer
Improving voice transfer data.
Training speech translation models with voice transfer typically amounts to supervised training on synthetic paired sequences of the same speaker. In particular, CVSS-T (Jia et al., 2022b) is the standard training set for S2ST with voice transfer and provides such artificial targets. However, Figure 4 shows that the average speaker similarity—as measured by the cosine similarity between speaker embeddings of source and target— on this dataset is very low with an average of $0.23$. As a calibration, state-of-the-art cross-lingual voice transfer lies around $0.40$ (Rubenstein et al., 2023). We thus also regenerate CVSS-T with our alignment-aware TTS, as it allows for voice transfer. As shown in Figure 4, the resynthesized CVSS-T displays a higher similarity, with an average of $0.47$. Yet, our training mixture which combines synthetic data and resynthesized CVSS-T still covers a wide range, with a significant mass below $0.40$.
Conditional training.
[…]
Classifier-free guidance.
Following Kreuk et al. (2023) we can increase the impact of the conditioning by using classifier-free guidance. We compute logits both with conditionings   very_good and   very_bad and sample from:
$\gamma l_{t,q}^{\leavevmode\nobreak\ \texttt{very\_good}}+(1-\gamma)l_{t,q}^{\leavevmode\nobreak\ \texttt{very\_bad}},$ (7)
which is compatible with real-time inference by producing both sets of logits with a batch size of 2. Section 4.6 shows that it significantly improves voice transfer.
4 Experiments
4.1 Architectural hyper-parameters
Hibiki consists of a Temporal Transformer with a latent dimension of 2560 (7040 for the SiLU gating), 24 layers, 20 heads and local attention over 500 tokens, i.e. 2.2B parameters and a 40s context. The Depth Transformer initially follows  Défossez et al. (2024), i.e. 6 layers per codebook, a latent dimension of 1024 (2816 for the gating), and 64 dimensions per head. It models $Q=16$ audio codebooks for the output stream, and the same for the input stream (only at training). To reduce the footprint of the Depth Transformer we distill it post-training into a smaller one, with 4 layers per codebook, weight sharing for codebooks 9 to 16, low-rank embedding tables of dimension 128, and a dimension of 2048 for the gating. This reduces its size from 1.1B parameters to 449M parameters, for a total of 2.7B parameters. We also train Hibiki-M as a distilled version of Hibiki, a 1.7B variant capable of running real-time on device, with a Temporal Transformer with a latent space of dimension 2048, and 16 layers, and only 8 codebooks levels per stream.
4.2 Training protocol
We train a French-English speech translation system through the following steps, each with a cosine learning rate schedule and AdamW (Loshchilov & Hutter, 2019), with a weight decay of 0.1, and momentum parameters of (0.9, 0.95).
Text pretraining. We first pretrain the Temporal Transformer from scratch on multilingual text-only data using next token prediction, for 600K steps, with a batch of 1,024 sequences of length 4,096. We use a cosine learning rate schedule, with 2K warmup steps and a maximum value of $4.8\cdot 10^{-4}$. Our training dataset is made of filtered web pages from Common Crawl, as well as curated sources such as Wikipedia, StackExchange or scientific articles and it contains 12.5% of multilingual documents.
Audio pretraining. Starting from the pretrained text model, we perform an audio pretraining on non-parallel French and English data with a single stream as done by Défossez et al. (2024). We train for 1,450K steps with a batch size of 144 and a learning rate of $2\cdot 10^{-4}$. Then we duplicate the weights of the Depth Transformer for multistream modeling.
Speech translation training. We build a French-English speech translation dataset of approximately 40K hours in each language. Starting from a collection of expressive audio content in French, we extract roughly 2.3M single speaker utterances each with a duration around 60 seconds. We transcribe these segments with Whisper (Radford et al., 2023), using the large-v3 model. We rely on PySBD (Sadvilkar & Neumann, 2020) to segment each transcript into sentences and use MADLAD-3B (Kudugunta et al., 2023) to translate them individually, before joining them back into a translated English transcript. We synthesize each with the TTS described in Section 3.2.2, conditioned on the original French speaker identity with a 10s utterance. We apply the silence insertion technique described in Section 3.2.2 to obtain simultaneous interpretation pairs. We train for 150K steps with a batch size of 96, a learning rate of $3\cdot 10^{-5}$ and compute the loss on both the source and the target streams. We use conditional training on speaker similarity as explained in 3.3 and apply noise augmentation techniques on the source audio. For each training pair, we introduce a special input EOS token on all audio tokens of the source at the first frame after the end of its speech utterance and use another special EOS token on the text tokens stream to indicate the end of the model speech utterance.
Speech translation fine-tuning. We use alignment-aware TTS generations introduced in Section 3.2.2 to build a synthetic dataset composed of long-form utterances and an improved version of CVSS-T/train, with natural pauses and high speaker similarity, totaling close to 900 hours. We fine-tune for 8K steps with a batch size of 8, a learning rate of $2\cdot 10^{-6}$, conditional training on the speaker similarity, special EOS tokens, and apply the loss to both streams.
Training of Hibiki-M. It goes through the same text and audio pre-training stages. During the speech translation training it is soft distilled from Hibiki, before going through the same fine-tuning step (without distillation).
Table 1: Comparison with offline baselines. We also report performance from a closed-source streaming model (*) as it uses the same evaluation protocol.
Model ASR-BLEU
Translatotron (Jia et al., 2019) 17.0
Translatotron 2 (Jia et al., 2022a) 26.0
S2UT (Lee et al., 2022a) 22.2
UnitY (Inaguma et al., 2023) 27.8
DASpeech (Fang et al., 2023) 25.0
RNN-Transducer* (Zhao et al., 2024) 25.4
StreamSpeech (Offline) (Zhang et al., 2024a) 28.5
Hibiki 30.5
Table 2: Objective comparison of Hibiki with StreamSpeech (Zhang et al., 2024a) and Seamless (Barrault et al., 2023).
Short-form (CVSS-C Fr-En test) Long-form (Audio-NTREX)
ASR Speaker End ASR Speaker End
Model BLEU ($\uparrow$) BLEU ($\uparrow$) Sim. ($\uparrow$) Offset ($\downarrow$) LAAL ($\downarrow$) BLEU ($\uparrow$) BLEU ($\uparrow$) Sim. ($\uparrow$) Offset ($\downarrow$) LAAL ($\downarrow$)
StreamSpeech 26.4 25.4 - 1.6 2.8 0.1 0.1 - N/A N/A
Seamless 37.0 33.8 0.30 1.4 2.8 25.4 23.9 0.43 1.6 4.2
Hibiki-M 37.5 33.7 0.34 2.8 3.5 25.9 25.0 0.39 2.3 5.5
Hibiki 39.2 35.5 0.41 2.9 3.4 27.5 26.6 0.48 2.7 5.0
4.3 Evaluation metrics and baselines
We first compare Hibiki to several offline baselines as well as a closed source streaming baseline (Zhao et al., 2024). We then perform a comparison between Hibiki and the two existing methods for simultaneous translation: Seamless (Barrault et al., 2023) and StreamSpeech (Zhang et al., 2024a) with a chunk size of 2560ms.
Translation quality. We evaluate translation quality by transcribing generated speech and computing a BLEU score (Post, 2018) with the reference, referred to as ASR-BLEU. When comparing to offline baselines in Table 1, we replicate the setting of Zhang et al. (2024a) for a fair comparison. However, we observe that this ASR model makes frequent mistakes, so in subsequent experiments we rather transcribe using Whisper medium  (Radford et al., 2023) and evaluate the BLEU score between ground-truth and hypothesis after normalization.333github.com/openai/whisper/blob/main/whisper/normalizers Since Seamless, StreamSpeech and Hibiki also produce a text translation, we also evaluate their BLEU score using the same text normalization.
Audio quality and naturalness. Human raters evaluate the audio quality of generated speech and its naturalness. We evaluate the latter as a proxy for “realism”: are the flow and prosody natural, are pauses smooth and properly placed or are there abrupt cuts? We compute each score for each model by averaging Mean Opinion Scores between 1 and 5 across 30 samples, each sample being evaluated by 15 raters.
Cross-lingual speaker similarity.
For objective evaluation, we use a standard model for speaker verification444github.com/microsoft/UniSpeech (“WavLM Large”) based on WavLM (Chen et al., 2022) and report the cosine similarity between the embeddings of the source and the generated speech. To mitigate potential biases due to using the same speaker verification model for conditional training (see Section 3.3), we also collect human judgments where raters are asked to rate the similarity to the source audio.
Latency. A metric for S2ST latency is the End Offset, which is the time (in seconds) between the end of the last word of the source and that of the last word in the output. We also measure the Length-Adaptative Average Lagging (LAAL) following the method described by Papi et al. (2022): it approximates the average time (in seconds) between the pronunciation of a source word and its translation, without requiring word-level alignments. We rely on word-level emission timestamps $(d_{i})_{1\dots n_{\mathrm{gen}}}$ produced by Whisper for $n_{\mathrm{gen}}$ words in the generated speech. We define $\delta=\frac{\Delta_{\mathrm{source}}}{max(n_{\mathrm{gen}},n_{\mathrm{ref}})}$ where $\Delta_{\mathrm{source}}$ is the duration of the source speech and $n_{\mathrm{ref}}$ the number of words in the reference translation. The LAAL score is then computed as $\frac{1}{n_{\mathrm{max}}}\sum_{i=1}^{n_{\mathrm{max}}}d_{i}-(i-1)\delta$ where $n_{max}=\mathrm{min}\{i|d_{i}\geq\Delta_{\mathrm{source}}\}$.
4.4 Evaluation datasets
Short-form data. We evaluate models on the Fr-En task of CVSS (Jia et al., 2022b). While it is the standard benchmark for S2ST and allows comparisons with previous models, we observe that 99% of its sequences are shorter than 10 seconds. We thus extend our evaluation to long-forms.
Long-form data. We collect long-form speech translations by recording bilingual speakers as they read Fr-En translations from the NTREX (Aepli et al., 2023) text corpus. This speech corpus, that we name Audio-NTREX contains 10 hours of real human speech in each language with 10 different speakers and an average of 50 sec. per utterance.
Real interpretation. To compare with human interpreters, we use 90 real interpretations of the European Parliament from VoxPopuli (Wang et al., 2021a) where translations contain the source speech at a lower volume. For a fair comparison, we also add the lowered source speech to generations of Hibiki and Seamless (see our external webpage for samples).
4.5 Inference configuration
We encode audio with the streaming codec and feed the tokens to Hibiki while decoding the output tokens to obtain streaming translation. At the end of the input, we send the EOS token to our model, and keep sampling until it produces its own EOS. Inference parameters are cross validated independently for each dataset using a held-out 8% of Audio-NTREX and the valid split of CVSS-C. The optimal parameters are $\gamma=3.0$, a temperature of 0.8, top-k of 250 for audio tokens and 50 for text tokens for Audio-NTREX. On CVSS, the same configuration is used except for text tokens that are sampled with a temperature of 0.1. We conjecture that the lower text temperature typically improves translation but can lead to producing an EOS token too early.
4.6 Results
Translation quality.
Table 1 compares Hibiki with offline baselines that have access to the complete source audio when translating. Despite performing simultaneous translation, Hibiki outperforms all models, including the offline variant of StreamSpeech. Table 2 benchmarks Hibiki against available baselines for simultaneous translation. In the short-form setting, our model outperforms StreamSpeech and Seamless at the cost of an average 0.7s of additional lagging. The long-form dataset represents a more significant challenge, as StreamSpeech does not manage to produce intelligible translations. Hibiki outperforms Seamless, again with a latency higher by an average of 0.8s.
Table 3: Human evaluation. Raters report Mean Opinion Scores (MOS) between 1 and 5.
Model Quality Speaker Sim. Naturalness
Ground-truth 4.18 $\pm$ 0.07 - 4.12 $\pm$ 0.08
Seamless 2.22 $\pm$ 0.08 2.86 $\pm$ 0.12 2.18 $\pm$ 0.09
Hibiki 3.78 $\pm$ 0.09 3.43 $\pm$ 0.10 3.73 $\pm$ 0.09
Audio fidelity.
Objective evaluations for speaker similarity, as reported in Table 2, show that Hibiki demonstrates significantly better voice transfer than Seamless (we do not evaluate StreamSpeech as it does not perform voice transfer). Human evaluations reported in Table 3 confirm this result and furthermore show a much higher quality and naturalness than Seamless, that get close to that of ground-truth audio from professional human interpreters. This means that not only Hibiki produces high-quality audio, but that it inserts smooth and natural pauses into its flow.
Table 4: Ablations.
Model ASR-BLEU LAAL
No lag 4.2 2.46
Constant lag (2s) 10.0 2.49
Constant lag (10s) 22.5 9.02
Sentence alignment 25.6 21.49
No Inner Monologue 17.1 14.34
No audio pretraining 14.6 5.12
Hibiki 26.6 5.0
Table 5: Ablations on classifier-free guidance.
CFG param. ASR-BLEU Speaker Sim.
No CFG 26.0 0.42
$\gamma=3.0$ (default) 26.6 0.48
$\gamma=10.0$ 18.9 0.44
Figure 5: Batched inference speed of Hibiki on a H100 SXM.
Ablation: alignment strategies. We compare our contextual alignment method with alternatives. Table 4 shows that applying no lag to the target speech during training results in very low translation quality, which can be expected since the model lacks context to produce the translation. Adding lag in training examples improves the ASR-BLEU, with 10 seconds representing a reasonable value, however the resulting average latency (as represented by LAAL) is much worse than using a contextual alignment, as the model does not adapt its flow to the context. A middle-ground between a constant lag and contextual alignment is that of “sentence alignment” which simply moves the start of each output sentence to the end of the corresponding source sentence. This improves translation quality, however degrading the latency even more. Overall, contextual alignment provides the best trade-off between translation quality and latency.
Ablation: Classifier-free guidance. Table 5 shows that using the very_good label provides a speaker similarity of 0.42, similar to that of Seamless (0.43). Using classifier-free guidance with $\gamma=3.0$ significantly improves it without significantly hurting translation quality, while increasing its weight too much results in degraded performance due to unintelligible speech. Supplementary material interestingly illustrates how increasing $\gamma$ to extreme values results in an exaggerated French accent (the source language in our experiments), which we can attribute to biases in the speaker model used to label our data.
Ablation: General ablations. Section 3.1.3 describes how jointly predicting text tokens serves as a scaffolding for audio generation. Table 4 illustrates this claim as training Hibiki in a unimodal fashion, without predicting text outputs, results much worse performance, as does starting from a pretrained text LM and training directly for S2ST.
Figure 6: Inference speed of Hibiki-M on an iPhone 16 Pro.
4.6.1 Inference capabilities
Batched inference. Hibiki’s inference uses temperature sampling with a constant framerate. This makes it easy to perform streaming classifier-free guidance as well as batching the processing of several sources of speech at the same time. This is unlike Seamless and StreamSpeech whose complex inference policies are difficult to batch as they require dynamic, irregular decisions for each sequence. Figure 5 shows that Hibiki remains faster than real-time on a H100 even when processing 320 sequences in parallel (or 160 with classifier-free guidance).
On-device inference. Our distilled Hibiki-M is competitive with Seamless on short-form and long-form translation as shown in Table 2. We attribute the lower speaker similarity on long-form audio to the lower number of quantizers modeled by Hibiki-M (8 instead of 16) which results in a twice lower audio bitrate. Figure 6 shows inference traces of Hibiki-M on an iPhone 16 Pro. Hibiki-M remains faster than real-time along a minute of inference, even with a batch size of 2 which is necessary for classifier-free guidance. Training Hibiki-M with sliding window attention would furthermore improve real-time factor along time.
4.6.2 Limitations
This study focuses on a single translation task (French to English) and scaling to more languages could benefit from MADLAD which is massively multilingual, however it would require training TTS systems on more languages. Moreover, while Hibiki reaches 35.5 ASR-BLEU against CVSS-C ground-truth targets, it reaches 47.9 ASR-BLEU if compared to MADLAD text translations instead. This shows that Hibiki is excellent at predicting translations that could be produced by MADLAD, and training it to predict pseudo-targets from better or more diverse translations can improve translation quality w.r.t ground-truth targets.
5 Conclusion
We introduce Hibiki, a model for simultaneous speech and text translation. Hibiki leverages a multistream architecture that casts live translation as simple temperature sampling, thanks to a weakly-supervised method for aligning paired data. Hibiki is competitive with the state-of-the-art in terms of translation quality, while demonstrating a much more natural flow and better voice transfer, close to human interpretation. Moreover, Hibiki is compatible with streaming batched inference, which facilitates large-scale deployment, while the smaller Hibiki-M runs in real-time on a smartphone.
Appendix A Appendix
Figure 7: Example of contextual alignments. We compute the contextual alignment ($a_{j}^{\text{ctx}}$) for four different samples and plot the associated input timestamps. Some results as the two at the bottom present extreme spikes meaning output words referring to input words very far in the future. These spikes are considered as anomalies and are smoothed out as explained in Section 3.2.2.